Find 2 consecutive positive odd intergers whose product is 143.
Answer by mananth(16946)

let the numbers be x, and x+2
x(x+2)=143
x^2+2x-143=0
x^2+13x-11x-143=0
x(x+13)-11(x+13)=0
(x+13)(x-11)=0
x= -13 or 11

The numbers can be 11 & 13 or -13, -11